Grade 11Magical Mathematics[Interesting Approach]

If f(x)+f(x+1/2) =1

Find limit 0 to 2 integration of f(x)dx.

ans is 1..plzz explain.

Ranjita yadav
13 Years agoGrade 11

2 Answers

raj razi
ApprovedApproved Tutor Answer13 Years ago

f(x)+f(x+1/2) =1

f(x)=1-f(x+1/2)

02f(x)dx=021-f(x+1/2)dx

02f(x)dx=2-02f(x+1/2)dx

take (x+1/2)=v

dx=dv

02f(v)dv=2-02f(v)dv

2(02f(v)dv)=2

02f(v)dv=1

02f(x)dx=1


Ranjita yadav
13 Years ago

But when u took x+1/2 as v , the limitss will be changed